OpenStudy (anonymous):

Given that r(t)•r’(t)=0 for all values of t, prove that ||r|| = c (constant) Given that r(t)•r’(t)=0 for all values of t, prove that ||r|| = c (constant) @Mathematics

OpenStudy (anonymous):

r(t)•r'(t) = aa'+bb'+cc'

OpenStudy (anonymous):

Not sure how that proves ||r|| = c must be constant though

OpenStudy (jamesj):

Consider the function f(t) = r(t)•r(t). Then f'(t) = r'(t)•r(t) + r(t)•r'(t) = 2 r(t)•r'(t) Now if r(t)•r'(t) = 0, then f'(t) = 0 therefore .... can you fill in the rest now?

OpenStudy (anonymous):

hmm

OpenStudy (jamesj):

Remember that \[ ||r||^2 = r . r \]

OpenStudy (anonymous):

Ah yea

OpenStudy (anonymous):

Ok, that explains it, just gotta think it over a bit. thanks :)

OpenStudy (jamesj):

We know now that \( f(t) = constant \). And by definition \( f(t) = ||r||^2 \) Hence ...

OpenStudy (anonymous):

Is letting f(t) = r(t)•r(t) the only way to do it?

OpenStudy (anonymous):

Hence ||r||= f(t)^2 = c^2

OpenStudy (jamesj):

square roots

OpenStudy (anonymous):

square roots?

OpenStudy (jamesj):

Not squared. f(t) = ||r||^2 = c implies ||r|| = sqrt(c) not c^2

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

But if f(t) = c

OpenStudy (anonymous):

and f(t) = ||r||^2

OpenStudy (anonymous):

ok, it should be f(t) = ||r||^2 = c?

OpenStudy (jamesj):

Yes

OpenStudy (jamesj):

And therefore \( ||r|| = \sqrt{c} \), a constant.

OpenStudy (anonymous):

I wouldnt have thought to set f(t) = r(t)•r(t), is that the only way to prove this?

OpenStudy (jamesj):

I just introduced f(t) to make it a bit easier to handle. But you do the same thing this way: \[ (||r||^2)' = (r.r)' = r.r' + r'.r = 2r.r' = 0 \ \ \implies \ \ ||r||^2 = c \ \ \implies \ ||r|| = \sqrt{c} \]

OpenStudy (jamesj):

As to whether there is another way to prove it, nothing comes to mind right now. Nonetheless, this is a short and elegant proof; why not use it?

OpenStudy (anonymous):

Yea, thanks for the help

OpenStudy (anonymous):

I am having trouble with this step: (r.r)′=r.r′+r′.r

OpenStudy (anonymous):

is that r * r or r dot r?

OpenStudy (anonymous):

I see the product rule being used so thats whats making me think it is r times r

OpenStudy (jamesj):

definitely the dot product between r and r: r . r. The expression r * r, r times r, is meaningless as these are vectors.

OpenStudy (anonymous):

so the derivative of the dot product is the same as the derivative of a product?

OpenStudy (anonymous):

r.r′+r′.r cause isnt that the product rule?

OpenStudy (jamesj):

It's a lemma/basic result in vector calculus that if u(t) and v(t) are vector functions, that the product rule holds: \[ \frac{d \ }{dt} u(t) . v(t) = u'(t) . v(t) + u(t) . v'(t) \]

OpenStudy (jamesj):

You should prove that for yourself by writing u and v in components, multiplying it out and differentiating.

OpenStudy (anonymous):

hmm, ok, I wont take anymore of your time with this problem, ill analyze what you've written for a better understanding, thanks :)

OpenStudy (jamesj):

ok